Exam 3: Probability

arrow
  • Select Tags
search iconSearch Question
flashcardsStudy Flashcards
  • Select Tags

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: Consider a sample space defined by events A1,A2,B1,B2.Let P(A1)= 0.40 ,P(B1 ∣ A1)= 0.60 and P(B1 ∣ A2)= 0.70 -What is P(A1 ∣ B1)?

(Essay)
4.8/5
(86)

If the firm interviewed was expecting medium or high growth,what is the probability of the firm being located in the West?

(Multiple Choice)
5.0/5
(77)

An office of six people is plagued by high absenteeism.It is thought that the probability that an employee is absent on a particular day is 0.03.Assuming that the event that one person is absent on a particular day is independent of the absence of any other employee,what is the probability that at least one employee is absent tomorrow?

(Multiple Choice)
4.9/5
(78)

If the union of several events covers the entire sample space,it is said these events are:

(Multiple Choice)
4.9/5
(82)

Which of the following is true?

(Multiple Choice)
4.8/5
(84)

If two events are mutually exclusive,they must also be collectively exhaustive.

(True/False)
4.9/5
(74)

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: Consider a sample space defined by events A1,A2,B1,B2.Let P(A1)= 0.40 ,P(B1 ∣ A1)= 0.60 and P(B1 ∣ A2)= 0.70 -What is P(A1 ∩ B2)?

(Essay)
4.8/5
(78)

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: Consider a sample space defined by events A1,A2,B1,B2.Let P(A1)= 0.40 ,P(B1 ∣ A1)= 0.60 and P(B1 ∣ A2)= 0.70 -What is P(A1 ∣ B2)?

(Essay)
4.8/5
(73)

Marginal probability is the probability that a given event will occur,with no other events taken into consideration.

(True/False)
4.8/5
(82)

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: A company wishes to evaluate the effectiveness of a marketing campaign.Seventy five percent of all potential professors were reached in a focused advertising program.Twenty eight percent of those contacted adopted the book while 8% of the adoptions came from professors who did not receive the promotional material.Define the following events of interest: A1 = Professor received advertising material A2 = Professor did not receive advertising material B1 = Professor adopts the book B2 = Professor does not adopt the book -What is the probability that a professor did not receive advertising material and does not adopt the book?

(Essay)
4.9/5
(76)

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: A table of joint probabilities is shown below,regarding the instructors at the University of Michigan,where the events Ai(i = 1,2,3)and Bj (j = 1,2)are defined as follows: A1 = Full professor,A2 = Associate professor,A3 = Assistant professor,B1 = Male instructor,and B2 = Female instructor. THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: A table of joint probabilities is shown below,regarding the instructors at the University of Michigan,where the events A<sub>i</sub>(i = 1,2,3)and B<sub>j</sub> (j = 1,2)are defined as follows: A<sub>1</sub> = Full professor,A<sub>2</sub> = Associate professor,A<sub>3</sub> = Assistant professor,B<sub>1</sub> = Male instructor,and B<sub>2</sub> = Female instructor.    -Show that P(A<sub>1</sub> ∣ B<sub>1</sub>)+ P(A<sub>2</sub> ∣ B<sub>1</sub>)+ P(A<sub>3</sub> ∣ B<sub>1</sub>)= 1 -Show that P(A1 ∣ B1)+ P(A2 ∣ B1)+ P(A3 ∣ B1)= 1

(Essay)
4.7/5
(66)

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: Consider a sample space defined by events A1,A2,B1,B2.Let P(A1)= 0.40 ,P(B1 ∣ A1)= 0.60 and P(B1 ∣ A2)= 0.70 -What is P(A2 ∣ B2)?

(Essay)
4.7/5
(76)

The posterior probability of any event must be at least as large as its prior probability.

(True/False)
4.8/5
(81)

Express the probability P(A |B)in words.

(Essay)
4.7/5
(76)

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: A table of joint probabilities is shown below,regarding the instructors at the University of Michigan,where the events Ai(i = 1,2,3)and Bj (j = 1,2)are defined as follows: A1 = Full professor,A2 = Associate professor,A3 = Assistant professor,B1 = Male instructor,and B2 = Female instructor. THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: A table of joint probabilities is shown below,regarding the instructors at the University of Michigan,where the events A<sub>i</sub>(i = 1,2,3)and B<sub>j</sub> (j = 1,2)are defined as follows: A<sub>1</sub> = Full professor,A<sub>2</sub> = Associate professor,A<sub>3</sub> = Assistant professor,B<sub>1</sub> = Male instructor,and B<sub>2</sub> = Female instructor.    -Calculate the marginal probabilities of events B<sub>j</sub>. -Calculate the marginal probabilities of events Bj.

(Essay)
4.9/5
(85)

Which of the following statements is true given that the events A and B have nonzero probabilities?

(Multiple Choice)
4.9/5
(69)

The type of diagram that shows subsets of the population on different branches is called a:

(Multiple Choice)
4.8/5
(58)

An example of the classical approach to probability would be:

(Multiple Choice)
4.8/5
(74)

Which of the following is not an example of the relative frequency approach to probability?

(Multiple Choice)
4.8/5
(80)

THE NEXT QUESTIONS ARE BASED ON THE FOLLOWING INFORMATION: A recent survey showed that 15% of computer programmers have experienced some form of wrist pain from typing,and that 25% are taking aspirin daily.Six percent of all programmers have both experienced some form of wrist pain from typing and taken aspirin on a daily basis. -What is the probability that a programmer has wrist pain and does not take aspirin on a daily basis?

(Multiple Choice)
4.9/5
(85)
Showing 221 - 240 of 345
close modal

Filters

  • Essay(0)
  • Multiple Choice(0)
  • Short Answer(0)
  • True False(0)
  • Matching(0)